Shane borrowed $1000 on a 5.6% discounted loan for a period of 20 months. A. What is the loan's discount. B. What is the net amount of money Shane receives? C. What is the loan's actual rate of interest? (To the nearest hundredth of a percent)
Answer by mananth(16946)


Shane borrowed $1000 on a 5.6% discounted loan for a period of 20 months.
I = P * r * t
I = 1000 * 0.056 *(20/12) =============== 5.6 /100 = 0.056
I = $93.33
Net amount = 1000-93.33 =906.7
I=P*r*t
93.33 =906.7 *r*(20/12)
93.3 *12/(20 *906.7) = r
r=0.0617
= 6.17 %